OpenStudy (anonymous):

Solve this system of equations using the addition method. 3x-y=9 2x+y=6 Solve this system of equations using the addition method. 3x-y=9 2x+y=6 @Mathematics

OpenStudy (anonymous):

5x + 0 = 15 x = 3 plug to the first equation 2.3 + y = 6 y = 0

OpenStudy (anonymous):

so i would just add the numbers then divide

OpenStudy (anonymous):

first add the equation 1 and 2 so that it will get rid of one of the variable In this example, you got it easy because in equation 1 you have -y, and in equation 2 you have y, so the sum of these equations will get rid of y

OpenStudy (anonymous):

in case you want to get rid of x first. then you would have to do the following: multiply 1st equation by 2 multiply 2nd equation by -3 so that it would appear like this 6x .... -6x ------------- 0x ......

OpenStudy (anonymous):

im sorry im very confused..see im takin algebra online and dnt have a teacher so i have to take tests online and have no idea how to do them

OpenStudy (anonymous):

ok so you have two equations and two variables which is x, y. They ask you to find those values of x, y so that when you plug them in to any equation, they have to be true. In this case if you plug x = 0 and y = 3. Both equations hold. That's how it goes anyways.

OpenStudy (anonymous):

Now they asked you to use addition method which means. some how you have to add equation 1 and equation 2

OpenStudy (anonymous):

3x-y=9 + 2x+y=6 ---------- 5x + 0 = 15 the new equation only has x now. So you know what x is, just plug into any of the equations they gave you, you will get y value.

OpenStudy (anonymous):

give me another problem, i'll explain in more details because this is problem is very simple.

OpenStudy (anonymous):

ok 4a-5b=7 4a+5b=17

OpenStudy (anonymous):

alright. this is also simple. Let's change to 4a + 5b = 7 2a + 3b = 9 instead.

OpenStudy (anonymous):

now what do you do? I want to know how you do it, then I'll explain

OpenStudy (anonymous):

i add 4a to 2a right

OpenStudy (anonymous):

continue

OpenStudy (anonymous):

then i add5b and3b

OpenStudy (anonymous):

ok. but the new equation doesn't get rid of any variables right? so your stuck

OpenStudy (anonymous):

you have to modify the original equations somehow so when adding you will get either 0x or 0y

OpenStudy (anonymous):

look at this problem you gave me: 4a-5b=7 4a+5b=17 notice the -5b (1st) and 5b (2nd equation). when adding the result is 0b. that's why i said they gave you a simple problem. Just adding.

OpenStudy (anonymous):

did you learn how to modify the equations yet? or just purely adding stuff?

OpenStudy (anonymous):

no did learn to modify the equations

OpenStudy (anonymous):

okay then please try to modify this one: 4a + 5b = 7 2a + 3b = 9

OpenStudy (anonymous):

i cnt i dnt understand how to modify im sorry

OpenStudy (anonymous):

let's try to modify so when adding I will get 0a ok?

OpenStudy (anonymous):

i see first equation have 4a. the 2nd have 2a. ... I need to have some how to have 2nd equation to have -4a. Right?

OpenStudy (anonymous):

ummm yes

OpenStudy (anonymous):

you can do this by multiply 2nd equation by -2

OpenStudy (anonymous):

what's the modified equation now?

OpenStudy (anonymous):

ok the modified (2nd equation) becomes: -4a -6b = -18

OpenStudy (anonymous):

now you can add the first and the modified.

OpenStudy (anonymous):

so then id add the 4a and 5b

OpenStudy (anonymous):

no no. it's gonna be like this: 4a + 5b = 7 + -4a -6b = -18 ----------------------- 0a -b = -11

OpenStudy (anonymous):

you only add the a and a... b and b not a and b.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so now you can solve for the b value. Once you have one variable solved, plug to any of the original equation, then bam. That's it

OpenStudy (anonymous):

can you modify to get rid of b instead of a this time? it's a bit harder, but same method. 4a + 5b = 7 2a + 3b = 9

OpenStudy (anonymous):

you have to modify both equations for b.

OpenStudy (anonymous):

4a+2a right

OpenStudy (anonymous):

yes what about b

OpenStudy (anonymous):

5b+3b?

OpenStudy (anonymous):

so you will get 6a + 8b = 16?

OpenStudy (anonymous):

it doesn't get rid of b does it?

OpenStudy (anonymous):

so i subtract

OpenStudy (anonymous):

you need to get help from professor about this. But those problems you gave me. you can just add them directly without modifying anything.

OpenStudy (anonymous):

you are not clear about the concept of adding.

OpenStudy (anonymous):

one thing I can tell you that when adding you must get rid of one of the variables.

OpenStudy (anonymous):

no im not im sorry thnks for tryin to help tho

OpenStudy (anonymous):

i have to go home. sorry